# AMix-1: A Pathway to Test-Time Scalable Protein Foundation Model

### v1

@@ 5.2 In Silico: AMix-1 Excels at In-Context Learning
The input consists of a randomly sampled protein family from CAMEO. As shown in Figure 7(a), AMix-1 generates novel proteins whose predicted 3D structures remain highly consistent with the input, as measured by structural alignment metric TM-score. The average predicted pLDDT exceeds 85, indicating confident and foldable outputs. Despite sequence-level novelty (low Seq-Sim to any input MSA sequence), the generated structures preserve the canonical fold, highlighting the model’s capacity to generalize conserved scaffolds.
@@
In Figure 7(b), we examine a challenging scenario where the input is an orphan protein—i.e., one with only a few known homologs and minimal representation in existing protein families. Such proteins are underrepresented in training data, making them difficult for models to generalize to. Despite this, AMix-1 successfully generates foldable variants that exhibit coherent tertiary structures with high pLDDT scores. This demonstrates that even sparse sequence profiles can serve as effective prompts, and that AMix-1 is capable of extrapolating structural patterns purely from weak but positionally localized evolutionary signals.
@@
We next assess functional generalization. Figure 7(c) presents results from prompting with an MSA containing enzymes annotated with EC 4.2.3.120 (beta-pinene synthase). The generated proteins are evaluated using CLEAN, a state-of-the-art model for EC number prediction. Over 80% of the generations are assigned the correct EC number with high confidence, and the predicted functional entropy remains low, indicating strong specificity. The average predicted CLEAN confidence for generated sequences closely matches the input enzyme, demonstrating functional conservation through in-context prompting.
@@ 5.3 Wet Lab: AMix-1 Designs High-Activity Proteins
Using our AMix-1-650M model, we generated 40 candidate sequences, each with ten or fewer mutations. The fitness of these variants was quantified using a fluorescent reporter assay, where higher DNA binding affinity results in stronger repression of a fluorescent protein, measured as “fold repression”. A higher fold repression value indicates superior function. The performance of AMix-1 was benchmarked against three baselines: (1) Wild Type (WT); (2) a Single-Mutant Library; and (3) EvoAI [57], a state-of-the-art directed evolution method. Illustration of wet lab verification process guided iteratively by AMix-1 is shown in Figure 8 while detailed protocols are provided in Section 14.
@@ Results.
The full experimental results are summarized in Figure 9. The superior performance of the AMix-1-designed variants highlights our model’s key strengths. It demonstrates a powerful ability to navigate the vast sequence-function landscape, identifying not just individual high impact mutations but also beneficial combinatorial effects. This capability allows AMix-1 to discover highly active sequences inaccessible to methods that rely on local exploration or the recombination of known variants.
@@ 6.1 Scaling AMix-1 with Evolutionary Algorithm
Operationally, we instantiate EvoAMix-1 as an evolutionary algorithm guided by a verifier that estimates fitness, inspired by the principles of natural selection (see Figure 10). Given a wild-type MSA, we encode it as a profile, a categorical distribution over amino acids, and use AMix-1 to generate a diverse batch of candidate variants. Leveraging its in-context learning capability in Section 5, AMix-1 produces variants that inherently preserve the structural and functional characteristics of the input. Each variant is then evaluated by a verifier that assigns a fitness score. Owing to AMix-1 ’s stochastic nature, the generated sequences span a range of fitness levels, with some surpassing all previous variants. High-fitness candidates are retained, while low-fitness ones are discarded. The surviving variants are used to construct a pseudo-MSA, which is then compressed into a new profile. This profile serves as the input to the next round of variant generation. By iteratively applying this propose-verify-update cycle, EvoAMix-1 progressively refines its outputs, producing increasingly fit protein sequences in a test-time scalable fashion. A detailed algorithm is provided in Algorithm 3.
@@ 6.2 A Proposal Distribution Perspective
We now turn to a more formal discussion of the mechanism underlying EvoAMix-1 , framing it through the lens of proposal distributions as shown in Figure 11. This perspective not only clarifies the optimization dynamics of our method, but also unifies traditional evolutionary algorithms in computational biology under a common framework, highlighting key distinctions between those approaches and EvoAMix-1 .
@@ 6.3 AMix-1 Exhibits Robust Test-Time Scaling across Diverse Tasks
We applied EvoAMix-1 to a series of protein design tasks. The results in Figure 12 show that it consistently exhibits strong test-time scaling behavior across all tasks—namely, as the number of verifier calls increases, the model continues to produce progressively higher-quality designs. Specifically, we evaluate EvoAMix-1 across three representative classes of protein design tasks:
@@
We report performance with respect to the verification budget, i.e., the number of verifier calls, as this directly reflects wet-lab throughput in real-world protein design scenarios. This perspective offers practical insight for lab-in-the-loop frameworks. More implementation details of these experiments can be found in Section 12. We now proceed to a more detailed discussion of the experimental results.
@@
To better understand the contribution of each component in EvoAMix-1 , we conduct targeted ablation experiments presented in Section 13. These studies reveal that removing either the pretrained generative model or the iterative profile refinement mechanism leads to substantial drops in both reward optimization and biological plausibility. Notably, substituting AMix-1 with random mutation causes the model to quickly drift away from structurally and functionally coherent solutions, while disabling prompt updates prevents convergence to high-fitness regions. These findings validate the importance of valid generation and adaptive conditioning in realizing effective test-time scaling.
@@ 7 Conclusion
Chaochen Gao9, Jingjing Gong2,3, Yuxuan Song2,3, Xiaoqing Zheng5, Deyi Xiong6, Lei Bai1, Ya-Qin Zhang3, Wei-Ying Ma3, Bowen Zhou1,4

### v3

@@ 5.2 In Silico: AMix-1 Excels at In-Context Learning
The input consists of a randomly sampled protein family from CAMEO. As shown in Figure˜7(a), AMix-1 generates novel proteins whose predicted 3D structures remain highly consistent with the input, as measured by structural alignment metric TM-score. The average predicted pLDDT exceeds 85, indicating confident and foldable outputs. Despite sequence-level novelty (low Seq-Sim to any input MSA sequence), the generated structures preserve the canonical fold, highlighting the model’s capacity to generalize conserved scaffolds.
@@
In Figure˜7(b), we examine a challenging scenario where the input is an orphan protein—i.e., one with only a few known homologs and minimal representation in existing protein families. Such proteins are underrepresented in training data, making them difficult for models to generalize to. Despite this, AMix-1 successfully generates foldable variants that exhibit coherent tertiary structures with high pLDDT scores. This demonstrates that even sparse sequence profiles can serve as effective prompts, and that AMix-1 is capable of extrapolating structural patterns purely from weak but positionally localized evolutionary signals.
@@
We next assess functional generalization. Figure˜7(c) presents results from prompting with an MSA containing enzymes annotated with EC 4.2.3.120 (beta-pinene synthase). The generated proteins are evaluated using CLEAN, a state-of-the-art model for EC number prediction. Over 80% of the generations are assigned the correct EC number with high confidence, and the predicted functional entropy remains low, indicating strong specificity. The average predicted CLEAN confidence for generated sequences closely matches the input enzyme, demonstrating functional conservation through in-context prompting.
@@ 5.3 Wet Lab: AMix-1 Designs High-Activity Proteins
Using our AMix-1-650M model, we generated 40 candidate sequences, each with ten or fewer mutations. The fitness of these variants was quantified using a fluorescent reporter assay, where higher DNA binding affinity results in stronger repression of a fluorescent protein, measured as “fold repression”. A higher fold repression value indicates superior function. The performance of AMix-1 was benchmarked against three baselines: (1) Wild Type (WT); (2) a Single-Mutant Library; and (3) EvoAI [57], a state-of-the-art directed evolution method. Illustration of wet lab verification process guided iteratively by AMix-1 is shown in Figure˜8 while detailed protocols are provided in Section˜14.
@@ Results.
The full experimental results are summarized in Figure˜9. The superior performance of the AMix-1-designed variants highlights our model’s key strengths. It demonstrates a powerful ability to navigate the vast sequence-function landscape, identifying not just individual high impact mutations but also beneficial combinatorial effects. This capability allows AMix-1 to discover highly active sequences inaccessible to methods that rely on local exploration or the recombination of known variants.
@@ 6.1 Scaling AMix-1 with Evolutionary Algorithm
Operationally, we instantiate EvoAMix-1 as an evolutionary algorithm guided by a verifier that estimates fitness, inspired by the principles of natural selection (see Figure˜10). Given a wild-type MSA, we encode it as a profile, a categorical distribution over amino acids, and use AMix-1 to generate a diverse batch of candidate variants. Leveraging its in-context learning capability in Section˜5, AMix-1 produces variants that inherently preserve the structural and functional characteristics of the input. Each variant is then evaluated by a verifier that assigns a fitness score. Owing to AMix-1 ’s stochastic nature, the generated sequences span a range of fitness levels, with some surpassing all previous variants. High-fitness candidates are retained, while low-fitness ones are discarded. The surviving variants are used to construct a pseudo-MSA, which is then compressed into a new profile. This profile serves as the input to the next round of variant generation. By iteratively applying this propose-verify-update cycle, EvoAMix-1 progressively refines its outputs, producing increasingly fit protein sequences in a test-time scalable fashion. A detailed algorithm is provided in Algorithm˜3.
@@ 6.2 A Proposal Distribution Perspective
We now turn to a more formal discussion of the mechanism underlying EvoAMix-1 , framing it through the lens of proposal distributions as shown in Figure˜11. This perspective not only clarifies the optimization dynamics of our method, but also unifies traditional evolutionary algorithms in computational biology under a common framework, highlighting key distinctions between those approaches and EvoAMix-1 .
@@ 6.3 AMix-1 Exhibits Robust Test-Time Scaling across Diverse Tasks
We applied EvoAMix-1 to a series of protein design tasks. The results in Figure˜12 show that it consistently exhibits strong test-time scaling behavior across all tasks—namely, as the number of verifier calls increases, the model continues to produce progressively higher-quality designs. Specifically, we evaluate EvoAMix-1 across three representative classes of protein design tasks:
@@
We report performance with respect to the verification budget, i.e., the number of verifier calls, as this directly reflects wet-lab throughput in real-world protein design scenarios. This perspective offers practical insight for lab-in-the-loop frameworks. More implementation details of these experiments can be found in Section˜12. We now proceed to a more detailed discussion of the experimental results.
@@
To better understand the contribution of each component in EvoAMix-1 , we conduct targeted ablation experiments presented in Section˜13. These studies reveal that removing either the pretrained generative model or the iterative profile refinement mechanism leads to substantial drops in both reward optimization and biological plausibility. Notably, substituting AMix-1 with random mutation causes the model to quickly drift away from structurally and functionally coherent solutions, while disabling prompt updates prevents convergence to high-fitness regions. These findings validate the importance of valid generation and adaptive conditioning in realizing effective test-time scaling.
@@ 7 Conclusion
Chaochen Gao9, Jingjing Gong2,3, Yuxuan Song2,3, Xiaoqing Zheng5, Deyi Xiong6, Lei Bai1, Wanli Ouyang1, Ya-Qin Zhang3, Wei-Ying Ma3,10, Bowen Zhou1,4